# Mass Translation and Reflectivity Device Solar Sail Momentum Management Using Predictive Control

### v2

@@ 3.5.1 Formulation of Iterative Backwards-in-Time MPC with Quantization
$\mbf{x}_{k+1}$ $=e^{\mbf{A}_{\ell}(t_{k+1}-t_{k})}\mbf{x}_{k}+\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{w}(t_{k})\mbf{w}(\tau)\textrm{d}\tau+\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{u1}(t_{k})\mbf{u}^{\text{AMT}}(\tau)\textrm{d}\tau$
$\hskip 20.0pt+\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{u2}(t_{k})u^{\text{RCD}}(\tau)\textrm{d}\tau$ (22)
@@
$\mbf{x}_{k+1}=\mbf{A}_{k}\mbf{x}_{k}+\mbf{B}_{w,k}^{-}\mbf{w}_{k}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{k}\\ \mbf{u}^{\text{AMT}}_{k+1}\end{bmatrix}+\mbf{A}_{c}\mbf{B}_{u2,c}^{-}u^{\text{RCD}}_{\text{max}},$
@@
$\mbf{x}_{j+1|t_{k}}$ $=\mbf{A}_{k}\mbf{x}_{j|t_{k}}+\mbf{B}_{w,k}^{-}\mbf{w}_{j|t_{k}}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{j|t_{k}}\\ \mbf{u}^{\text{AMT}}_{j+1|t_{k}}\end{bmatrix}+\mbf{B}_{u2,k}^{-}u^{\text{RCD}}_{j|t_{k}},\hskip 5.74997pt\text{for}\hskip 5.74997ptj=0,1,\cdots,N-N_{bk}-1,$
$\mbf{x}_{j+1|t_{k}}$ $=\mbf{A}_{k}\mbf{x}_{j|t_{k}}+\mbf{B}_{w,k}^{-}\mbf{w}_{j|t_{k}}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{j|t_{k}}\\ \mbf{u}^{\text{AMT}}_{j+1|t_{k}}\end{bmatrix}+\mbf{A}_{c}\mbf{B}_{u2,c}^{-}u^{\text{RCD}}_{\text{max}},\hskip 5.74997pt\text{for}\hskip 5.74997ptj=N-N_{bk},\cdots,N-1,$
@@
where $N_{bk}=1,2,\cdots,N-1$ is the number of iterations performed since the initial MPC solution and is also the number of timesteps at the end of the horizon with a quantized RCD input.
@@
The proposed iterative backwards-in-time MPC policy operates by performing a sequential PWM quantization to the furthest optimal RCD input in the prediction horizon, and iteratively re-optimize the remainder of the unmodified sequence with $N_{bk}$ fixed PWM-quantized RCD inputs at the end of the prediction horizon. Note that every iteration of the iterative backwards MPC algorithm removes one more time step of RCD input from the end of the prediction horizon with a fixed PWM-based quantization value of $\mbf{A}_{c}\mbf{B}_{u2,c}^{-}u^{\text{RCD}}_{\text{max}}$. In the first iteration, only the RCD input from the final time step in the prediction horizon is fixed as a PWM on-off pulse, while the AMT inputs remain free for optimization. In subsequent iterations, one additional RCD input is quantized and fixed, progressing step by step closer to the current time step. The modified dynamics are incorporated into the MPC policy as equality constraints, replacing the continuous-time model with the PWM-based quantized inputs as appropriate. The iterative process continues $N_{bk}$ times until $N_{bk}=N-1$, at which point all RCD inputs in the prediction horizon are quantized other than the first time step. Finally, the first time step’s input is extracted and implemented as a PWM-quantized input, as outlined in Section 3.3. This method ensures that the MPC solution incorporates knowledge of PWM quantization, improving dynamic model fidelity in the MPC prediction. Fig. 5 illustrates the sequential iterative backward MPC algorithm, showing how the far-end RCD input is fixed at each iteration and the remaining design variables are re-optimized. Since fixed quantized RCD inputs no longer contribute to the cost function, their influence is gradually removed during iterations of the backwards iteration procedure. To maintain consistent weighting in the cost function, a scaling factor of $N/(N-N_{bk})$ is applied to the remaining RCD input design variables. This ensures that the optimization problem remains consistent across iterations, yielding an optimal solution that respects the quantization constraints.
@@ 3.5.2 Summary of MPC Policy for Momentum Management Strategy 2
$\operatorname*{minimize}_{{\bm{\mathcal{X}}},\hskip 2.0pt{\bm{\mathcal{U}}},\hskip 2.0pt{\bm{\alpha}}}\sum_{j=0}^{N-1}\bigg{(}\mbf{x}_{j|t_{k}}^{\top}\mbf{Q}\mbf{x}_{j|t_{k}}+\mbf{u}_{j|t_{k}}^{\text{AMT}^{\top}}\mbf{R}\mbf{u}_{j|t_{k}}^{\text{AMT}}+\tilde{\mbf{u}}^{\text{AMT}^{\top}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}+\frac{N}{N-N_{bk}}\mbf{u}_{j|t_{k}}^{\text{RCD}^{\top}}\mbf{R}\mbf{u}_{j|t_{k}}^{\text{RCD}}\bigg{)}$
$\hskip 260.0004pt+\mbf{x}_{N|t_{k}}^{\top}\mbf{P}\mbf{x}_{N|t_{k}}+\mbf{u}_{N|t_{k}}^{\top}\mbf{R}_{N}\mbf{u}_{N|t_{k}}+{\bm{\alpha}}^{\top}\mbf{C}{\bm{\alpha}}$ (23)
@@
$\hskip 20.00003pt\mbf{x}_{j+1|t_{k}}=\begin{cases}&\mbf{A}_{k}\mbf{x}_{j|t_{k}}+\mbf{B}_{w,k}^{-}\mbf{w}_{j|t_{k}}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{j|t_{k}}\\ \mbf{u}^{\text{AMT}}_{j+1|t_{k}}\end{bmatrix}+\mbf{B}_{u2,k}^{-}u^{\text{RCD}}_{j|t_{k}},\hskip 5.74997pt\text{for}\hskip 5.74997ptj=0,1,\cdots,N-N_{bk}-1,\\ &\mbf{A}_{k}\mbf{x}_{j|t_{k}}+\mbf{B}_{w,k}^{-}\mbf{w}_{j|t_{k}}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{j|t_{k}}\\ \mbf{u}^{\text{AMT}}_{j+1|t_{k}}\end{bmatrix}+\mbf{A}_{c}\mbf{B}_{u2,c}^{-}u^{\text{RCD}}_{\text{max}},\hskip 5.74997pt\text{for}\hskip 5.74997ptj=N-N_{bk},\cdots,N-1,\end{cases}$
@@ 4 Numerical Simulation Results
A numerical simulation demonstrating the tracking performance of RWs PID control law is presented in Section 4.2, where the PID gains are tuned to track the desired attitude under a non-equilibrium initial attitude without any momentum management actuation. The PID attitude control law successively tracks the spacecraft attitude while the RWs angular momentum keeps growing under the consistent affect of disturbances. Section 4.3 presents results with the state-of-the-art momentum management strategy used on Solar Cruiser. Section 4.4 shows the simulation results of the proposed MPC momentum management strategies formulated in Sections 3.4 and 3.5.
@@ 4.1 Problem Setup
AMT range, $\mbf{u}^{\text{AMT}}_{\text{max}}=\Big{[}r^{\text{AMT}}_{b1,\text{max}}\,\,\,r^{\text{AMT}}_{b2,\text{max}}\Big{]}^{\top}$ $\big{[}\pm 0.29\,\,\,\pm 0.29\big{]}^{\top}$ m
AMT speed, $\dot{\mbf{u}}^{\text{AMT}}_{\text{max}}=\Big{[}\dot{r}^{\text{AMT}}_{b1,\text{max}}\,\,\,\dot{r}^{\text{AMT}}_{b2,\text{max}}\Big{]}^{\top}$ $\big{[}\pm 0.5\,\,\,\pm 0.5\big{]}^{\top}$ mm/s
@@
constant SRP force, $\mbf{f}^{\text{SRP}}_{b}$ $\big{[}0\,\,\,0\,\,\,0.013\big{]}^{\top}$ N
@@ 4.2 Attitude Tracking with RW PID Control
Numerical simulation are carried using the nonlinear solar sail attitude dynamics in (6). Given that the maximum translation rate of the AMT is $\pm 0.5$ mm/s for Solar Cruiser [4], it is assumed that $\ddot{\mbf{r}}_{b}^{ps}=\mbf{0}$ in the simulation, and a first-order hold on the AMT input allows for the substitution $\dot{\mbf{r}}_{b}^{ps}(t_{k})=\frac{\mbf{r}_{b}^{ps}(t_{k+1})-\mbf{r}_{b}^{ps}(t_{k})}{t_{k+1}-t_{k}}$ between time steps $t_{k},t_{k+1})$. The simulation results for the slew over $3000$ seconds without any momentum management is shown in Fig. 6. Fig. 6(a) illustrates the attitude tracking performance, while Fig. 6(b) depicts the evolution of angular momentum in the RWs. As the RWs adjust the solar sail’s attitude from ${\bm{\theta}}_{0}=\Big{[}2\,\,\,2\,\,\,1\Big{]}^{\top}$ deg to ${\bm{\theta}}_{d}=\mbf{0}$ deg and maintain the desired orientation, they simultaneously counteract the disturbance torque, leading to a steady accumulation of angular momentum. This buildup continues unabated, as shown in Fig. 6(b), where the RW angular momentum exceeds its maximum capacity (indicated by the red dashed line). In practice, this would result in RW saturation, rendering them ineffective for further attitude control. Although the disturbance in the roll-axis ($\underrightarrow{b}^{3}$ axis) is relatively small, the angular momentum still exhibits a consistent growth trend at steady state after the slew maneuver, underscoring the cumulative nature of angular momentum under constant disturbances.
@@
Figure 6: Simulation results of the spacecraft attitude and RWs angular momentum using a PID attitude control law without any momentum management control. The red dashed lines in Fig. 6(b) indicate the saturation limit of RWs.
@@ 4.3 Solar Cruiser’s PID Momentum Management
$u^{\text{AMT}}_{1}$ $=K^{\text{AMT}}_{p}h^{\text{RWs}}_{b2}+K^{\text{AMT}}_{d}\dot{h}^{\text{RWs}}_{b2}+K^{\text{AMT}}_{i}\int^{t}_{t_{0}}h^{\text{RWs}}_{b2}(\tau)\textrm{d}\tau,$
$u^{\text{AMT}}_{2}$ $=-K^{\text{AMT}}_{p}h^{\text{RWs}}_{b1}-K^{\text{AMT}}_{d}\dot{h}^{\text{RWs}}_{b1}-K^{\text{AMT}}_{i}\int^{t}_{t_{0}}h^{\text{RWs}}_{b1}(\tau)\textrm{d}\tau.$
@@
Simulation results with this momentum management strategy and the same initial conditions and parameters used in Section 4.2 are included in Fig. 7. The momentum management performance is similar to that in [11, 16], although this is difficult to compare quantitatively due to redacted plot axes in [11, 16]. The black dashed lines in Fig. 7(b) indicate the activation thresholds, while the green dashed lines represent the deactivation thresholds. The momentum management method from [11, 16] is effective at keeping the angular momentum of the RWs within reasonable bounds with realistic actuation inputs. However, it is worth noting that tuning the PID gains to obtain a satisfactory result required a significant amount of tuning.

### v3

@@ 4.5.1 Formulation of Iterative Backwards-in-Time MPC with Quantization
$\mbf{x}_{k+1}=e^{\mbf{A}_{\ell}(t_{k+1}-t_{k})}\mbf{x}_{k}+\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{w}(t_{k})\mbf{w}(\tau)\textrm{d}\tau\\ \\ +\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{u1}(t_{k})\mbf{u}^{\text{AMT}}(\tau)\textrm{d}\tau+\int^{t_{k+1}}_{t_{k}}e^{\mbf{A}_{\ell}(t_{k+1}-\tau)}\mbf{B}_{u2}(t_{k})u^{\text{RCD}}(\tau)\textrm{d}\tau$ (40)
@@
$\mbf{x}_{k+1}=\mbf{A}_{k}\mbf{x}_{k}+\mbf{B}_{w,k}^{-}\mbf{w}_{k}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{k}\\ \mbf{u}^{\text{AMT}}_{k+1}\end{bmatrix}+\mbf{A}_{c}\mbf{B}_{u2,c}^{-}u^{\text{RCD}}_{\text{max}},$ (41)
@@
$\mbf{x}_{j+1|t_{k}}=\mbf{A}_{k}\mbf{x}_{j|t_{k}}+\mbf{B}_{w,k}^{-}\mbf{w}_{j|t_{k}}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{j|t_{k}}\\ \mbf{u}^{\text{AMT}}_{j+1|t_{k}}\end{bmatrix}+\mbf{B}_{u2,k}^{-}u^{\text{RCD}}_{j|t_{k}},$ (42)
for $j=0,1,\cdots,N-N_{bk}-1$ and
$\mbf{x}_{j+1|t_{k}}=\mbf{A}_{k}\mbf{x}_{j|t_{k}}+\mbf{B}_{w,k}^{-}\mbf{w}_{j|t_{k}}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{j|t_{k}}\\ \mbf{u}^{\text{AMT}}_{j+1|t_{k}}\end{bmatrix}+\mbf{A}_{c}\mbf{B}_{u2,c}^{-}u^{\text{RCD}}_{\text{max}},$ (43)
@@
for $j=N-N_{bk},\cdots,N-1$, where $N_{bk}=1,2,\cdots,N-1$ is the number of iterations performed since the initial MPC solution and is also the number of timesteps at the end of the horizon with a quantized RCD input.
@@
The proposed iterative backwards-in-time MPC policy operates by performing a sequential PWM quantization to the furthest optimal RCD input in the prediction horizon, and iteratively re-optimize the remainder of the unmodified sequence with $N_{bk}$ fixed PWM-quantized RCD inputs at the end of the prediction horizon. Note that every iteration of the iterative backwards MPC algorithm removes one more time step of RCD input from the end of the prediction horizon with a fixed PWM-based quantization value of $\mbf{A}_{c}\mbf{B}_{u2,c}^{-}u^{\text{RCD}}_{\text{max}}$. In the first iteration, only the RCD input from the final time step in the prediction horizon is fixed as a PWM on-off pulse, while the AMT inputs remain free for optimization. In subsequent iterations, one additional RCD input is quantized and fixed, progressing step by step closer to the current time step. The modified dynamics are incorporated into the MPC policy as equality constraints, replacing the continuous-time model with the PWM-based quantized inputs as appropriate. The iterative process continues $N_{bk}$ times until $N_{bk}=N-1$, at which point all RCD inputs in the prediction horizon are quantized other than the first time step. Finally, the first time step’s input is extracted and implemented as a PWM-quantized input, as outlined in Section 4.3. This method ensures that the MPC solution incorporates knowledge of PWM quantization, improving dynamic model fidelity in the MPC prediction. Figure 5 illustrates the sequential iterative backward MPC algorithm, showing how the far-end RCD input is fixed at each iteration and the remaining design variables are re-optimized. Since fixed quantized RCD inputs no longer contribute to the cost function, their influence is gradually removed during iterations of the backwards iteration procedure. To maintain consistent weighting in the cost function, a scaling factor of $N/(N-N_{bk})$ is applied to the remaining RCD input design variables. This ensures that the optimization problem remains consistent across iterations, yielding an optimal solution that respects the quantization constraints.
@@ 4.5.2 Summary of MPC Policy for Momentum Management Strategy 2
$\operatorname*{minimize}_{{\bm{\mathcal{X}}},\hskip 2.0pt{\bm{\mathcal{U}}},\hskip 2.0pt{\bm{\alpha}}}\sum_{j=0}^{N-1}\bigg(\mbf{x}_{j|t_{k}}^{\top}\mbf{Q}\mbf{x}_{j|t_{k}}+\mbf{u}_{j|t_{k}}^{\text{AMT}^{\top}}\mbf{R}\mbf{u}_{j|t_{k}}^{\text{AMT}}+\tilde{\mbf{u}}^{\text{AMT}^{\top}}_{j|t_{k}}\tilde{\mbf{R}}\tilde{\mbf{u}}^{\text{AMT}}_{j|t_{k}}$
$\hskip 20.00003pt+\frac{N}{N-N_{bk}}\mbf{u}_{j|t_{k}}^{\text{RCD}^{\top}}\mbf{R}\mbf{u}_{j|t_{k}}^{\text{RCD}}\bigg)+\mbf{x}_{N|t_{k}}^{\top}\mbf{P}\mbf{x}_{N|t_{k}}+\mbf{u}_{N|t_{k}}^{\top}\mbf{R}_{N}\mbf{u}_{N|t_{k}}+{\bm{\alpha}}^{\top}\mbf{C}{\bm{\alpha}}$ (44)
@@
$\hskip 20.00003pt\mbf{x}_{j+1|t_{k}}=\begin{cases}&\mbf{A}_{k}\mbf{x}_{j|t_{k}}+\mbf{B}_{w,k}^{-}\mbf{w}_{j|t_{k}}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{j|t_{k}}\\ \mbf{u}^{\text{AMT}}_{j+1|t_{k}}\end{bmatrix}+\mbf{B}_{u2,k}^{-}u^{\text{RCD}}_{j|t_{k}},\\ &\hskip 137.99927pt\text{for}\hskip 5.74997ptj=0,1,\cdots,N-N_{bk}-1,\\ &\mbf{A}_{k}\mbf{x}_{j|t_{k}}+\mbf{B}_{w,k}^{-}\mbf{w}_{j|t_{k}}+\begin{bmatrix}\mbf{B}_{u1,k}^{-}&\mbf{B}_{u1,k}^{+}\end{bmatrix}\begin{bmatrix}\mbf{u}^{\text{AMT}}_{j|t_{k}}\\ \mbf{u}^{\text{AMT}}_{j+1|t_{k}}\end{bmatrix}+\mbf{A}_{c}\mbf{B}_{u2,c}^{-}u^{\text{RCD}}_{\text{max}},\\ &\hskip 137.99927pt\text{for}\hskip 5.74997ptj=N-N_{bk},\cdots,N-1,\end{cases}$
@@ 5 Numerical Simulation Results
A numerical simulation demonstrating the tracking performance of RWs PID control law is presented in Section 5.2, where the PID gains are tuned to track the desired attitude under a non-equilibrium initial attitude without any momentum management actuation. The PID attitude control law successively tracks the spacecraft attitude while the RWs angular momentum keeps growing under the consistent affect of disturbances. Section 5.3 presents results with the state-of-the-art momentum management strategy used on Solar Cruiser. Section 5.4 shows the simulation results of the proposed MPC momentum management strategies formulated in Sections 4.4 and 4.5.
@@ 5.1 Problem Setup
AMT range, $\Big[r^{\text{AMT}}_{b1,\text{max}}\,\,\,r^{\text{AMT}}_{b2,\text{max}}\Big]^{\top}$ $\big[\pm 0.29\,\,\,\pm 0.29\big]^{\top}$ m
AMT speed, $\Big[\dot{r}^{\text{AMT}}_{b1,\text{max}}\,\,\,\dot{r}^{\text{AMT}}_{b2,\text{max}}\Big]^{\top}$ $\big[\pm 0.5\,\,\,\pm 0.5\big]^{\top}$ mm/s
@@
constant SRP force, $\mbf{f}^{\text{SRP}}_{b}$ $\big[0.0003\,\,\,0\,\,\,0.013\big]^{\top}$ N
@@ 5.2 Attitude Tracking with RW PID Control
Numerical simulation are carried using the nonlinear solar sail attitude dynamics in (14). Given that the maximum translation rate of the AMT is $\pm 0.5$ mm/s for Solar Cruiser [10], it is assumed that $\ddot{\mbf{r}}_{b}^{ps}=\mbf{0}$ in the simulation, and a first-order hold on the AMT input allows for the substitution $\dot{\mbf{r}}_{b}^{ps}(t_{k})=\frac{\mbf{r}_{b}^{ps}(t_{k+1})-\mbf{r}_{b}^{ps}(t_{k})}{t_{k+1}-t_{k}}$ between time steps $t_{k},t_{k+1})$. The simulation results for the slew over $3000$ seconds without any momentum management is shown in Figure 6. Figure 6(a) illustrates the attitude tracking performance, while Figure 6(b) depicts the evolution of angular momentum in the RWs. As the RWs adjust the solar sail’s attitude from ${\bm{\theta}}_{0}=\Big[2\,\,\,0\,\,\,1\Big]^{\top}$ deg to ${\bm{\theta}}_{d}=\mbf{0}$ deg and maintain the desired orientation, they simultaneously counteract the disturbance torque, leading to a steady accumulation of angular momentum. This buildup continues unabated, as shown in Figure 6(b), where the RW angular momentum exceeds its maximum capacity (indicated by the red dashed line). In practice, this would result in RW saturation, rendering them ineffective for further attitude control. Although the disturbance in the roll-axis ($\underrightarrow{b}^{3}$ axis) is relatively small, the angular momentum still exhibits a consistent growth trend at steady state after the slew maneuver, underscoring the cumulative nature of angular momentum under constant disturbances.
@@
Figure 6: Simulation results of the spacecraft attitude and RWs angular momentum using a PID attitude control law without any momentum management control. The red dashed lines in Figure 6(b) indicate the saturation limit of RWs.
@@ 5.3 Solar Cruiser’s PID Momentum Management
$u^{\text{AMT}}_{1}$ $=K^{\text{AMT}}_{p}h^{\text{RWs}}_{b2}+K^{\text{AMT}}_{d}\dot{h}^{\text{RWs}}_{b2}+K^{\text{AMT}}_{i}\int^{t}_{t_{0}}h^{\text{RWs}}_{b2}(\tau)\textrm{d}\tau,$ (45)
$u^{\text{AMT}}_{2}$ $=-K^{\text{AMT}}_{p}h^{\text{RWs}}_{b1}-K^{\text{AMT}}_{d}\dot{h}^{\text{RWs}}_{b1}-K^{\text{AMT}}_{i}\int^{t}_{t_{0}}h^{\text{RWs}}_{b1}(\tau)\textrm{d}\tau.$ (46)
@@
Simulation results with this momentum management strategy and the same initial conditions and parameters used in Section 5.2 are included in Figure 7. The momentum management performance is similar to that in [20, 26], although this is difficult to compare quantitatively due to redacted plot axes in [20, 26]. The black dashed lines in Figure 7(b) indicate the activation thresholds, while the green dashed lines represent the deactivation thresholds. The momentum management method from [20, 26] is effective at keeping the angular momentum of the RWs within reasonable bounds with realistic actuation inputs. However, it is worth noting that tuning the PID gains to obtain a satisfactory result required a significant amount of tuning.